All questions please
Question 4 A particular securitys equilibrium rate of return is 10% For all securities, the inflation risk premium is 2.75 percent and the real interest rate is 3 percent. The securitys liquidity risk premium is 50 percent and maturity risk premium is .75 percent. The security has no special covenants. What is the securitys default risk premium? Hint:i IP+RIR+ DRP+LRP+SCP+ MRP 2% 4% 5% 6% Question 5 1 pts Suppose we observe the following rates: 1R1-08, 1R2 12, and E(2r1)- 14. If the liquidity premium theory of the term structure of interest rates holds, what is the liquidity premium for year 2 215% О 6.15% 845% О 12.45% 14.00%
Answer #1

SEE THE IMAGE. ANY DOUBTS, FEEL FREE TO ASK. THUMBS UP PLEASE

Home nert Page Layout Formulas Data Review View dd-Ins Σ AutoSum Cut ー E ゴWrap Text Sort &Find & Conditional Format CeInsert Delete Format Formatting as Table Styles2 Clear aCopy B า 프 . Ej-., Δ. : r_一 逻锂函Merge & Center. $, % , 弼,8 Format Painter Cells Edting Number Alignment Clipboard AB AC AD 213 214 215 216 217 218 219 220 221 i-= IP+ RIR + DRP + LRP + SCP + MRP 10%-2.75% + 3% + DRP +0.5% +0% + 0.75% 1096-796 + DRP DRP = 3% 223 224 225 226 227 228 229 230 231 WARRANT fund bond CLEAN INVOICE PRICE 14 EUAC Aw PM LIFE LP IP MRP. INFLATION YIELD - YIELD SPOT Sheet2 AFN BANKING NOTE BIDDING, UNDERWRITING 19-01-2019Home nert Page Layout Formulas Data Review View dd-Ins Cut Σ AutoSum Calibri ー E ゴWrap Text General ta copy. ,_a. ars-函Merge & Center, $, % , 弼,8 C Paste B 1 u. Conditional Format CeInsert Delete Format Formatting, as Table w styles. ▼ ㆆ ▼ Sort &Find & 2 ClearFe Select Edting Format Painter Clipboard EP110 EC Font Alignment Number Styles Cells ED EF EG EH El EJ EK EL EM EN EO EP EQ 93 94 95 96 97 98 ACCORDING TO LIQUIDITY PREMIUM THEORY OF THE TERM STRUCTURE OF INTEREST RATES 1R2-[(1+1R1)(1hE(2R1)+L2)]^1/2-1 0.12 [(1+0.08)(1+0.14+L2)]A1/2 -1 1.12 = [(1.08)( 1.14扎2)M/2 100 101 102 103 104 105 106 107 108 109 110 111ッ SQUARING BOTH THE SIDES 1.2544 (1.08) (1.14 L2)] 1.2544/1.08 1.14 L2 1.16148 = 1.14 + L2 0.02148 L2 2.1596-L2 ANSWER: L2-2.15% YIELD SPOT Sheet2 AFN BANKING NOTE BIDDING, UNDERWRITING EUAC AW PM LIFE LP I, MR INFLATION YIELD WARRANT fund bond CLEAN INVOICE PRICE 014 - 福 130% 06:31 19-01-2019